# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

### v2

@@ A.3 Implementation Details
Zero-Shot. Once pre-trained, our CAPE framework can be directly utilized for zero-shot forecasting where the model remains frozen and no parameter is updated. Similar to the MOMENT model Goswami et al. (2024), we retain the pre-trained reconstruction head and mask the last patch of the input to perform forecasting: $\hat{\mathbf{y}}=\hat{\mathbf{x}}_{[T-c:T]}$.
Model Details. We design our model by stacking 4 layers of the CAPE encoder, each with a hidden size of 512 and 4 attention heads. For environment representations, we incorporate 16 distinct environments, each encoded with a size of 512. To ensure a fair comparison, PatchTST is configured with the same number of layers and hidden size as our CAPE-based model. For all other baseline models, we adopt the architectures as reported in previous studies Wang et al. (2024a); Kamarthi & Prakash (2023); Panagopoulos et al. (2021).
Training Details. For the training process, we pre-train CAPE, PEM, and PatchTST on a single Nvidia A100 GPU. During pre-training, we utilize only 70% of the available training data, specifically the first 70% of the dataset for each disease category. We set the learning rate to $1\times 10^{-5}$ and apply a dropout rate of 0.1 to prevent overfitting. In the CAPE pre-training strategy, we assign a weight of 0.5 to $\alpha$ to balance the contributions of contrast loss to the whole loss function. A detailed illustration of our pre-training strategy is shown in Algorithm 1. After pre-training, we fine-tune the entire model using a single Nvidia K80 GPU, maintaining the same hyperparameter settings for consistency. The best-performing model is selected based on its performance on the validation set. Similarly, for all baseline models, we train each until convergence and select the optimal model based on validation set performance for the subsequent test.
@@
A.4 Full Results on Pre-train datasets
@@
A.5 Full results for few-shot forecasting
@@
A.6 Impact of pre-train ratio on the downstream datasets
We provide additional evaluations for CAPE on downstream datasets to analyze the impact of the pre-training ratio. As shown in Figure 9, increasing the pre-training ratio eventually improves downstream performance across all datasets.
@@ A.6 Impact of pre-train ratio on the downstream datasets
A.7 Visualization of the Estimated Environments
@@
A.8 Visualization of Distribution Shift for Downstream Datasets
We provide a visualization of the sample distribution used in this study. Each sample has a fixed length of 36, representing the historical infection trajectory. To better understand the distributional differences, we use t-SNE to reduce the data to one dimension and visualize the training and test samples using different colors. As shown in Figure 11, a significant distribution shift is visually apparent across most datasets. To quantitatively assess the distributional differences between the training and test sets, we calculate the Central Moment Discrepancy (CMD) score Zellinger et al. (2017). The CMD score measures the discrepancy between the central moments of the two distributions up to a specified order $K$. For two distributions $X$ (training set) and $X_{\text{test}}$ (test set), the CMD score is defined as:
$\quad\text{CMD}(X,X_{\text{test}})$ (25)
$=\|\mu_{1}(X)-\mu_{1}(X_{\text{test}})\|_{2}+\sum_{k=2}^{K}\|\mu_{k}(X)-\mu_{k}(X_{\text{test}})\|_{2},$
@@ A.8 Visualization of Distribution Shift for Downstream Datasets
A.9 Latent space visualization of Measle and Covid datasets from pre-trained models.
@@
A.10 Hyper-parameter Sensitivity Analysis
@@
$\mathcal{L}_{\text{final}}$ $=\sum_{\mathbf{x}\in\mathbf{X}}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+$ (27)
$\alpha\,\mathcal{L}_{\text{CL}}(\hat{\mathbf{E}}^{(L)},\hat{\mathbf{E}^{\prime}}^{(L)})+\frac{\beta}{L}\sum_{l}\mathcal{L}_{\text{CL}}(\tilde{{\mathbf{E}}}^{(l)},\tilde{{\mathbf{E}}^{\prime}}^{(l)}),$
$\mathbf{X}\sim\mathcal{D}^{\prime}_{s}\in\mathcal{D}_{\text{pre}}$

### v3

@@ A.5 Implementation Details
A.6 Evaluation Metrics
In this study, we primarily use MSE for its sensitivity to hard cases (large errors) and common use in epidemic forecasting  (Datilo et al., 2019). In addition, we also report MAPE and peak timing error (averaged over horizons 1–16) on the ILI USA dataset as an example:
Table 8: Model Performance Metrics
Model MAPE Peak Error
CAPE 0.188 1.226
Informer 0.246 1.553
LSTM 0.264 1.398
DLinear 0.225 1.442
A.7 Environment Influence
We also add Gaussian noise to environment weights (Equation 4), causing performance drops, which highlights the importance of accurate environment estimation:
Table 9: Model Performance under Different Noise Scales
Noise Scale (%) 5% 10% 15% 20%
MSE 0.327 0.373 0.456 0.559
MAPE 0.198 0.211 0.236 0.264
A.8 Full Results on Pre-train datasets
@@
A.9 Full results for few-shot forecasting
@@
A.10 Impact of the Environment Distribution and Pre-train Epochs
Figure 9: Downstream performance vs. environment distribution and pre-train epochs.
A.11 Impact of pre-train ratio on the downstream datasets
We provide additional evaluations for CAPE on downstream datasets to analyze the impact of the pre-training ratio. As shown in Figure LABEL:fig:pretrain_scale, increasing the pre-training ratio eventually improves downstream performance across all datasets.
@@ A.11 Impact of pre-train ratio on the downstream datasets
A.12 Mitigating Distribution Shift
Figure 11: CMD scores w/ and w/o environment estimate between train/test sets. We measure the CMD scores based on the last embedding output by our model.
A.13 Disentangling Disease-Specific Dynamics
After pre-training, both CAPE and PatchTST were frozen and applied to downstream datasets directly to generate latent space embeddings for each sample. We then employed the Davies-Bouldin Index (DBI) to assess the separability of these embeddings by disease. CAPE consistently achieved lower DBI scores (we show an example in Figure 16), demonstrating a superior ability to distinguish between different diseases in the latent space compared to PatchTST. This enhanced separability highlights the effectiveness of CAPE’s environment estimation and backdoor adjustment. These mechanisms are crucial for mitigating the confounding influence of noisy, spurious, and environment-dependent factors ($X_{sp}$), thereby enabling the model to better isolate and represent the unique, intrinsic, and causal disease dynamics ($X_{ca}$).
Figure 12: DBI between the embeddings of each pair of downstream datasets from the pre-trained model. (See Figure  16).
A.14 Visualization of the Estimated Environments
@@
A.15 Visualization of Distribution Shift for Downstream Datasets
We provide a visualization of the sample distribution used in this study. Each sample has a fixed length of 36, representing the historical infection trajectory. To better understand the distributional differences, we use t-SNE to reduce the data to one dimension and visualize the training and test samples using different colors. As shown in Figure 14, a significant distribution shift is visually apparent across most datasets. To quantitatively assess the distributional differences between the training and test sets, we calculate the Central Moment Discrepancy (CMD) score (Zellinger et al., 2017). The CMD score measures the discrepancy between the central moments of the two distributions up to a specified order $K$. For two distributions $X$ (training set) and $X_{\text{test}}$ (test set), the CMD score is defined as:
$\quad\text{CMD}(X,X_{\text{test}})=\|\mu_{1}(X)-\mu_{1}(X_{\text{test}})\|_{2}+\sum_{k=2}^{K}\|\mu_{k}(X)-\mu_{k}(X_{\text{test}})\|_{2},$ (21)
@@ A.15 Visualization of Distribution Shift for Downstream Datasets
A.16 Latent space visualization of Measle and Covid datasets from pre-trained models.
@@
A.17 Hyper-parameter Sensitivity Analysis
@@
$\mathcal{L}_{\text{final}}$ $=\sum_{\mathbf{x}\in\mathbf{X}}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+\alpha\,\mathcal{L}_{\text{CL}}(\hat{\mathbf{E}}^{(L)},\hat{\mathbf{E}^{\prime}}^{(L)})+\frac{\beta}{L}\sum_{l}\mathcal{L}_{\text{CL}}(\tilde{{\mathbf{E}}}^{(l)},\tilde{{\mathbf{E}}^{\prime}}^{(l)}),$ (23)
$\hskip 85.35826pt\mathbf{X}\sim\mathcal{D}^{\prime}_{s}\in\mathcal{D}_{\text{pre}}$